OpenStudy (anonymous):

The function f(x)=(x-4)/absolute(x-4), is defined everywhere except at x=4. If possible, define f(x) at 4 so that it becomes continuous at 4?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

uve gotta evaluate f(x) as x tends to 4

OpenStudy (anonymous):

a) Not possible because there is an infinite discontinuity at the given point. b) f (4) = 0 c) f (4) = -1 d) Not possible because there is a jump discontinuity at the given point. e) f (4) = 1

OpenStudy (anonymous):

\[\lim_{x \rightarrow 4+}[(x-4)/|x-4|] =\lim_{h \rightarrow 0}[(4+h-4)/|4+h-4|] = h/h = 1\]

OpenStudy (anonymous):

wait...no

OpenStudy (anonymous):

d

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

bcoz the LHL is -1

OpenStudy (anonymous):

for continuity function must be defined at that point, howz it 1?

OpenStudy (anonymous):

we have to evaluate the limit uzma...which is infact nonexistent..as LHL is -1 and RHL is 1

OpenStudy (anonymous):

we r to define f so that it is continuous at 4

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so akileez..see the left hand limit at 4 is -1, and the right hand limit is 1, so ur answer is d

OpenStudy (anonymous):

for x greater than 4, its 1 and for x less than 4 its -1,

OpenStudy (anonymous):

him is right now

OpenStudy (anonymous):

ok, I get it! since the left and right cannot equal, there is not point where f(x) will be continuous _ isn't his general for anything over absolutes?

OpenStudy (anonymous):

but, why is it a jump disc and not an infinite?

OpenStudy (anonymous):

because both the limit values are finite, theyre just not equal to each other

OpenStudy (anonymous):

and if they were equal to each other?

OpenStudy (anonymous):

for a function to be continuous L.H.L=R.H.L= value of the function at that point

OpenStudy (anonymous):

uzma can you pls write down the calcs that work out the value for the limit form the left and right side pls!